# Advancing Cross-domain Discriminability in Continual Learning of Vison-Language Models

### v1

@@ Abstract
Continual learning (CL) with Vision-Language Models (VLMs) has overcome the constraints of traditional CL, which only focuses on previously encountered classes. During the CL of VLMs, we need not only to prevent the catastrophic forgetting on incrementally learned knowledge but also to preserve the zero-shot ability of VLMs. However, existing methods require additional reference datasets to maintain such zero-shot ability and rely on domain-identity hints to classify images across different domains. In this study, we propose Regression-based Analytic Incremental Learning (RAIL), which utilizes a recursive ridge regression-based adapter to learn from a sequence of domains in a non-forgetting manner and decouple the cross-domain correlations by projecting features to a higher-dimensional space. Cooperating with a training-free fusion module, RAIL absolutely preserves the VLM’s zero-shot ability on unseen domains without any reference data. Additionally, we introduce Cross-domain Task-Agnostic Incremental Learning (X-TAIL) setting. In this setting, a CL learner is required to incrementally learn from multiple domains and classify test images from both seen and unseen domains without any domain-identity hint. We theoretically prove RAIL’s absolute memorization on incrementally learned domains. Experiment results affirm RAIL’s state-of-the-art performance in both X-TAIL and existing Multi-domain Task-Incremental Learning settings. The code will be released upon acceptance.
@@ 1 Introduction
Continual learning (CL) hadsell2020embracing ; de2021continual ; wang2024comprehensive is a crucial area in machine learning, which requires a learner to incrementally learn new data instead of training from scratch. The main challenge in CL is known as catastrophic forgetting mccloskey1989catastrophic , where learning new knowledge results in the forgetting of the old one. To this end, various CL approaches kirkpatrick2017overcoming ; li2017learning ; rebuffi2017icarl ; buzzega2020dark ; wang2024comprehensive have been proposed to solve the forgetting issue. As a typical CL setting, Class-Incremental Learning (CIL) (Fig. 1 (a)) aims to achieve robust discriminability on all seen classes. Despite the advancements, existing approaches mainly focus on classifying images only from seen classes, thereby limiting the model’s generalizability.
@@ 3.2 Datasets
In X-TAIL’s cross-domain setting, the learner should encompass as extensive data distributions as possible. Following previous work Zheng_2023_ICCV ; yu2024boosting ; gao2024clip ; radford2021learning ; Zhou_2022 , we select 10 different image-classification datasets from different domains for our setting: Aircraft maji2013finegrained , Caltech101 xu2016simple , DTD Saad_2020 , EuroSAT helber2019eurosat , Flowers flowers_dataset , Food bossard14 , MNIST 6296535 , OxfordPet oxfordpet , StanfordCars dehghan2017view , and SUN397 xiao2010sun . CL methods under X-TAIL should discriminate images from a total of $1,101$ classes across all domains.
@@ 3.3 Evaluation metrics
We adopt the evaluation metrics from Zheng_2023_ICCV for our setting. As illustrated in Fig. 2, each column represents the performance on a specific domain after each learning step, while the rows correspond to the learning sequence. In traditional CL settings, only the results in the lower diagonal where the learner has been exposed to the exemplars from the test domains are measured. Nevertheless, X-TAIL extends this evaluation to cover the entire matrix, recording performances across both learned and unlearned domains. The “Average” metric indicates the average accuracy of all learning steps across all domains. The “Last” metric, which is the average of performance across all domains after the final learning step, reflects the learner’s adaptability to new domains. Additionally, the average of the upper diagonal blocks, referred to as the “Transfer” metric, measures the extent to which the zero- shot ability is preserved throughout incremental learning.
@@ 4.1 Motivation
$\hat{\mathbf{y}}=\mathbf{X}_{e}\mathbf{W},\quad\operatorname*{arg\,min}_{\mathbf{W}}\left\|\mathbf{Y}-\mathbf{X}_{e}\right\|_{F}^{2}+\lambda\left\|\mathbf{W}\right\|_{F}^{2},$ (1)
@@
$\mathbf{\alpha}=\left(\mathbf{K}+\lambda\mathbf{I}\right)^{-1}\mathbf{Y},$ (3)
where $\mathbf{K}=\mathcal{K}\left(\mathbf{X},\mathbf{X}\right)$ denotes the covariance kernel matrix, and $\mathcal{K}\left(\cdot,\cdot\right)$ can be any positive-definite kernel function. The classification logits are derived as $\mathbf{\hat{y}}=\mathcal{K}\left(f_{I}\left(\mathbf{x}_{\text{test}}\right),\mathbf{X}\right)\mathbf{\alpha}$. Throughout the paper, we use the Radial Basis Function (RBF) kernel han2012parameter by default. The choice between primal and dual form ridge regression depends on whether the system is over-determined (more equations than unknowns) or under-determined (more unknowns than equations) gentle2012numerical . Details on the relationships between primal and dual ridge regression can be found in Appendix B.
@@ 4.2 Regression-based analytic incremental learning
Figure 5: RAIL Overview. (a) During inference, the fusion module utilizes the Zero-shot logits to identify whether a test image is aligned with seen or unseen classes. If classified as a seen class, the Fusion logits combine the RAIL-Adapter logits and the Zero-shot logits; otherwise solely rely on the Zero-shot logits. (b) Primal: at the $n$-th learning step, features $\mathbf{X}_{e}$ extracted by CLIP’s image encoder are projected to higher dimensional $\mathbf{\Phi}$ via RHL and then update the parameter $\mathbf{W}$ and memory $\mathbf{M}_{p}$ by Theorem 1. (c) Dual: features extracted by CLIP’s image encoder update the kernel $\mathbf{K}$, parameter $\mathbf{\alpha}$, and memory $\mathbf{M}_{d}$ by Theorem 2.
@@ 4.2.1 RAIL-Adapter
Similarly, the dual parameter $\mathbf{\alpha}^{\left(n\right)}$ satisfying Eqn. 4 can be obtained based on $\mathbf{\alpha}^{\left(n-1\right)}$, an updating kernel $\mathbf{K}^{\left(n\right)}$, and a memory matrix $\mathbf{M}_{d}^{\left(n\right)}$. We denote the matrix $\mathbf{C}^{\left(n\right)}$ as the concatenated one-hot label matrices of all $n$ domains. The solution is defined in Theorem 2.
@@
$\mathbf{\alpha}^{\left(n\right)}=\left(\mathbf{K}^{\left(n\right)}+\lambda\mathbf{I}\right)^{-1}\mathbf{C}^{\left(n\right)}$ (7)
@@
At each incremental learning step, the kernel matrix $\mathbf{K}$ updates recursively along the main diagonal, preserving the correlations among class-prototypes from all domains. During testing, the kernel covariance between the feature of test image extracted by CLIP and memory matrix $\mathbf{M}_{d}$ is calculated to obtain the classification logits $\hat{\mathbf{y}}=\mathcal{K}\left(f_{I}\left(\mathbf{x}_{\text{test}}\right),\mathbf{M}_{d}\right)\mathbf{\alpha}$. Specifically, $\mathbf{M}_{d}$ dynamically updates according to the data stream via concatenated class-prototypes. These class-prototypes can be the feature embeddings $\mathbf{X}_{e}$ extracted by CLIP, K-means centroids, or Gaussian Mixture Model means. We use raw feature embeddings $\mathbf{X}_{e}$ by default, which is sufficient to validate our method. The complete proofs for both theorems are provided in Appendix C.
@@ 5 Experiments
We evaluate RAIL method under both X-TAIL and MTIL settings, as mentioned in Sec. 3.2. The learning order is set alphabetically: Aircraft, Caltech101, DTD, EuroSAT, Flowers, Food, MNIST, OxfordPet, StanfordCars, and SUN397. Additional experiments with a random order are provided in Tab. 3 in Appendix H. To ensure compatibility with different domains, we follow the common practice of sampling a 16-shot training set for each domain, while using the original test set for evaluation gao2024clip ; zhang2021tip ; zhou2022conditional ; Zhou_2022 ; khattak2023maple . The implementation details are provided in Appendix D.
@@ 5.1 Comparison results
We follow the setting in yu2024boosting to evaluate our methods on the few-shot MTIL. In this context, the RAIL-Adapter reduces to domain-specific multi-head classifiers with provided domain-identity during testing. The comparison results are shown in Appendix G. RAIL consistently outperforms previous state-of-the-art approaches on all three metrics. These results demonstrate that the proposed non-linear projections significantly improve the separability of features extracted by CLIP. Consequently, the ridge regression-based classifier can effectively adapt the pre-trained model to new domains.
@@ Appendix A Algorithm details
         Obtain the dual parameter $\mathbf{\alpha}^{\left(1\right)}$ using Eqn. 7
@@
              Update the dual parameter $\mathbf{\alpha}^{\left(n\right)}$ using Eqn. 7
@@ Appendix B Connection between primal & dual ridge regression
where the former solution is based on the outer-product of data and the latter one is based on the inner-product of data. Parameter $\mathbf{W}$ can thus be rewritten as: $\mathbf{W}=\mathbf{\Phi}^{\top}\mathbf{\alpha}$ with $\mathbf{\alpha}=(\mathbf{\Phi}\mathbf{\Phi}^{\top}+\lambda\mathbf{I}_{n})^{-1}\mathbf{Y}$. In this way, the solution $\mathbf{W}$ is interpreted to lie in the span of the sample-cases, even if the dimensionality of the projected features $\mathbf{\Phi}$ is larger than the number of samples.
@@
$\mathbf{\hat{y}}=\phi\left(\mathbf{x}\right)\mathbf{W}=\phi\left(\mathbf{x}\right)\mathbf{\Phi}^{\top}(\mathbf{\Phi}\mathbf{\Phi}^{\top}+\lambda\mathbf{I}_{n})^{-1}\mathbf{Y}=\mathcal{K}\left(\mathbf{x},\mathbf{X}\right)\mathbf{\alpha},$ (12)
@@ Appendix C Proof of Theorems
$\mathbf{\alpha}^{\left(n\right)}=\left(\mathcal{K}\left(\mathbf{X}_{e}^{\left(1:n\right)},\mathbf{X}_{e}^{\left(1:n\right)}\right)+\lambda\mathbf{I}\right)^{-1}\mathbf{Y}^{\left(1:n\right)}.$ (21)
@@
The kernel matrix at $n$-th step for $\mathbf{\alpha}^{\left(n\right)}$ can be partitioned as:
@@
where the matrix $\mathbf{Y}^{\left(n\right)}$ consists of one-hot labels that are disjoint with those in previous $n-1$ domains. In this way, the parameter $\mathbf{\alpha}^{\left(n\right)}$ is solved by updating $\mathbf{K}^{\left(n\right)}$ and $\mathbf{C}^{\left(n\right)}$, resulting in an identical solution to the one of joint learning in Eqn. 21. Thus, the Theorem 2 is proved.

### v2

# Advancing Cross-domain Discriminability in Continual Learning of Vision-Language Models

@@ Abstract
Continual learning (CL) with Vision-Language Models (VLMs) has overcome the constraints of traditional CL, which only focuses on previously encountered classes. During the CL of VLMs, we need not only to prevent the catastrophic forgetting on incrementally learned knowledge but also to preserve the zero-shot ability of VLMs. However, existing methods require additional reference datasets to maintain such zero-shot ability and rely on domain-identity hints to classify images across different domains. In this study, we propose Regression-based Analytic Incremental Learning (RAIL), which utilizes a recursive ridge regression-based adapter to learn from a sequence of domains in a non-forgetting manner and decouple the cross-domain correlations by projecting features to a higher-dimensional space. Cooperating with a training-free fusion module, RAIL absolutely preserves the VLM’s zero-shot ability on unseen domains without any reference data. Additionally, we introduce Cross-domain Task-Agnostic Incremental Learning (X-TAIL) setting. In this setting, a CL learner is required to incrementally learn from multiple domains and classify test images from both seen and unseen domains without any domain-identity hint. We theoretically prove RAIL’s absolute memorization on incrementally learned domains. Experiment results affirm RAIL’s state-of-the-art performance in both X-TAIL and existing Multi-domain Task-Incremental Learning settings. The code is released at https://github.com/linghan1997/Regression-based-Analytic-Incremental-Learning.
@@ 1 Introduction
Continual learning (CL) hadsell2020embracing ; de2021continual ; wang2024comprehensive is a crucial area in machine learning, which requires a learner to incrementally learn new data instead of training from scratch. The main challenge in CL is known as catastrophic forgetting mccloskey1989catastrophic , where learning new knowledge results in the forgetting of the old one. To this end, various CL approaches kirkpatrick2017overcoming ; li2017learning ; rebuffi2017icarl ; buzzega2020dark ; wang2024comprehensive ; liu2024locality have been proposed to solve the forgetting issue. As a typical CL setting, Class-Incremental Learning (CIL) (Fig. 1 (a)) aims to achieve robust discriminability on all seen classes. Despite the advancements, existing approaches mainly focus on classifying images only from seen classes, thereby limiting the model’s generalizability.
@@ 3.2 Datasets
In X-TAIL’s cross-domain setting, the learner should encompass as extensive data distributions as possible. Following previous work Zheng_2023_ICCV ; yu2024boosting ; gao2024clip ; radford2021learning ; Zhou_2022 ; awt2024, we select 10 different image-classification datasets from different domains for our setting: Aircraft maji2013finegrained , Caltech101 xu2016simple , DTD Saad_2020 , EuroSAT helber2019eurosat , Flowers flowers_dataset , Food bossard14 , MNIST 6296535 , OxfordPet oxfordpet , StanfordCars dehghan2017view , and SUN397 xiao2010sun . Specifically, to prevent the redundancy of learning overlapping classes and to maintain the integrity of the setting, CIFAR-100 krizhevsky2009learning was excluded because it includes many classes that overlap with those in other domains. As a result, CL methods under X-TAIL should discriminate images from a total of $1,100$ classes across all domains.
@@ 3.3 Evaluation metrics
We adopt the evaluation metrics from Zheng_2023_ICCV for our setting. As illustrated in Fig. 2, each column represents the performance on a specific domain after each learning step, while the rows correspond to the learning sequence. In traditional CL settings, only the results in the lower diagonal where the learner has been exposed to the exemplars from the test domains are measured. Nevertheless, X-TAIL extends this evaluation to cover the entire matrix, recording performances across both learned and unlearned domains. The “Average” metric, averaged on the orange blocks, indicates the average accuracy of all learning steps across all domains. The gray and green blocks under the diagonal show the classification performance on these domains after the model has learned these domains. Specifically, the green blocks represent the model’s last performance on these domains after learning all domains. The “Last” metric, which is the average of the green blocks, reflects the learner’s adaptability to new domains. Additionally, the blue blocks in the upper-right matrix indicate the model’s zero-shot performance on these domains before learning these domains. The average of these blocks, referred to as the “Transfer” metric, measures the extent to which the zero-shot ability is preserved throughout incremental learning.
@@ 4.1 Motivation
$\hat{\mathbf{y}}=\mathbf{X}_{e}\mathbf{W},\quad\operatorname*{arg\,min}_{\mathbf{W}}\left\|\mathbf{Y}-\mathbf{X}_{e}\mathbf{W}\right\|_{F}^{2}+\lambda\left\|\mathbf{W}\right\|_{F}^{2},$ (1)
@@
$\boldsymbol{\alpha}=\left(\mathbf{K}+\lambda\mathbf{I}\right)^{-1}\mathbf{Y},$ (3)
where $\mathbf{K}=\mathcal{K}\left(\mathbf{X},\mathbf{X}\right)$ denotes the covariance kernel matrix, and $\mathcal{K}\left(\cdot,\cdot\right)$ can be any positive-definite kernel function. The classification logits are derived as $\mathbf{\hat{y}}=\mathcal{K}\left(f_{I}\left(\mathbf{x}_{\text{test}}\right),\mathbf{X}\right)\boldsymbol{\alpha}$. Throughout the paper, we use the Radial Basis Function (RBF) kernel han2012parameter by default. The choice between primal and dual form ridge regression depends on whether the system is over-determined (more equations than unknowns) or under-determined (more unknowns than equations) gentle2012numerical . Details on the relationships between primal and dual ridge regression can be found in Appendix B.
@@ 4.2 Regression-based analytic incremental learning
Figure 5: RAIL Overview. (a) During inference, the fusion module utilizes the Zero-shot logits to identify whether a test image is aligned with seen or unseen classes. If classified as a seen class, the Fusion logits combine the RAIL-Adapter logits and the Zero-shot logits; otherwise solely rely on the Zero-shot logits. (b) Primal: at the $n$-th learning step, features $\mathbf{X}_{e}$ extracted by CLIP’s image encoder are projected to higher dimensional $\mathbf{\Phi}$ via RHL and then update the parameter $\mathbf{W}$ and memory $\mathbf{M}_{p}$ by Theorem 1. (c) Dual: features extracted by CLIP’s image encoder update the kernel $\mathbf{K}$, parameter $\boldsymbol{\alpha}$, and memory $\mathbf{M}_{d}$ by Theorem 2.
@@ 4.2.1 RAIL-Adapter
Similarly, the dual parameter $\boldsymbol{\alpha}^{\left(n\right)}$ satisfying Eqn. 4 can be obtained based on $\boldsymbol{\alpha}^{\left(n-1\right)}$, an updating kernel $\mathbf{K}^{\left(n\right)}$, and a memory matrix $\mathbf{M}_{d}^{\left(n\right)}$. We denote the matrix $\mathbf{C}^{\left(n\right)}$ as the concatenated one-hot label matrices of all $n$ domains. The solution is defined in Theorem 2.
@@
$\boldsymbol{\alpha}^{\left(n\right)}=\left(\mathbf{K}^{\left(n\right)}+\lambda\mathbf{I}\right)^{-1}\mathbf{C}^{\left(n\right)}$ (7)
@@
At each incremental learning step, the kernel matrix $\mathbf{K}$ updates recursively along the main diagonal, preserving the correlations among class-prototypes from all domains. During testing, the kernel covariance between the feature of test image extracted by CLIP and memory matrix $\mathbf{M}_{d}$ is calculated to obtain the classification logits $\hat{\mathbf{y}}=\mathcal{K}\left(f_{I}\left(\mathbf{x}_{\text{test}}\right),\mathbf{M}_{d}\right)\boldsymbol{\alpha}$. Specifically, $\mathbf{M}_{d}$ dynamically updates according to the data stream via concatenated class-prototypes. These class-prototypes can be the feature embeddings $\mathbf{X}_{e}$ extracted by CLIP, K-means centroids, or Gaussian Mixture Model means. We use raw feature embeddings $\mathbf{X}_{e}$ by default, which is sufficient to validate our method. The complete proofs for both theorems are provided in Appendix C.
@@ 5 Experiments
We evaluate RAIL method under both X-TAIL and MTIL settings, as mentioned in Sec. 3.2. The learning order is set alphabetically: Aircraft, Caltech101, DTD, EuroSAT, Flowers, Food, MNIST, OxfordPet, StanfordCars, and SUN397. Additional experiments with a random order are provided in Tab. 3 in Appendix G. To ensure compatibility with different domains, we follow the common practice of sampling a 16-shot training set for each domain, while using the original test set for evaluation gao2024clip ; zhang2021tip ; zhou2022conditional ; Zhou_2022 ; khattak2023maple . The implementation details are provided in Appendix D.
@@ 5.1 Comparison results
We follow the setting in yu2024boosting to evaluate our methods on the few-shot MTIL, comparing it against the performance of baselines reported in yu2024boosting . In this context, RAIL is reduced to the structure of multiple domain-specific classifiers trained on each domain separately. The domain-identity guides the test image to the corresponding classifier for within-domain prediction. The comparison results are shown in Tab. 2. RAIL consistently outperforms previous state-of-the-art approaches on all three metrics. These results demonstrate that the proposed non-linear projections significantly improve the separability of features extracted by CLIP. Consequently, the ridge regression-based classifier can effectively adapt the pre-trained model to new domains.
Zero-shot 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 65.3
Fine-tune 30.6 93.5 76.8 65.1 91.7 92.9 83.3 96.6 84.9 65.4 71.3 77.5
Transfer
LwF 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
LwF-VR 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
WiSE-FT 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
ZSCL 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
MoE 87.9 68.2 44.1 48.1 64.7 88.8 69.0 89.1 64.5 65.1 68.9
Primal-RAIL 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
Dual-RAIL 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
Average
LwF 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
LwF-VR 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
WiSE-FT 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
ZSCL 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
MoE 30.0 89.6 73.9 58.7 69.3 79.3 88.1 76.5 89.1 65.3 65.8 71.4
Primal-RAIL 32.9 94.5 69.9 58.1 71.8 84.4 88.5 70.4 89.0 66.1 65.7 71.9
Dual-RAIL 36.0 94.2 70.9 58.8 70.6 84.3 88.5 70.3 89.7 66.5 65.8 72.3
Last
LwF 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
LwF-VR 22.9 89.9 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
WiSE-FT 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
ZSCL 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
MoE 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1
Primal-RAIL 32.9 95.1 70.3 63.2 81.5 95.6 88.5 89.7 89.0 72.5 71.0 77.2
Dual-RAIL 36.0 94.8 71.5 64.1 79.5 95.3 88.5 89.4 91.5 74.6 71.3 77.9
@@ 6 Conclusion
7 Acknowledgement
This research was supported by the National Natural Science Foundation of China (62306117) and the Guangzhou Basic and Applied Basic Research Foundation (2024A04J3681).
@@ Appendix A Algorithm details
         Obtain the dual parameter $\boldsymbol{\alpha}^{\left(1\right)}$ using Eqn. 7
@@
              Update the dual parameter $\boldsymbol{\alpha}^{\left(n\right)}$ using Eqn. 7
@@ Appendix B Connection between primal & dual ridge regression
where the former solution is based on the outer-product of data and the latter one is based on the inner-product of data. Parameter $\mathbf{W}$ can thus be rewritten as: $\mathbf{W}=\mathbf{\Phi}^{\top}\boldsymbol{\alpha}$ with $\boldsymbol{\alpha}=(\mathbf{\Phi}\mathbf{\Phi}^{\top}+\lambda\mathbf{I}_{n})^{-1}\mathbf{Y}$. In this way, the solution $\mathbf{W}$ is interpreted to lie in the span of the sample-cases, even if the dimensionality of the projected features $\mathbf{\Phi}$ is larger than the number of samples.
@@
$\mathbf{\hat{y}}=\phi\left(\mathbf{x}\right)\mathbf{W}=\phi\left(\mathbf{x}\right)\mathbf{\Phi}^{\top}(\mathbf{\Phi}\mathbf{\Phi}^{\top}+\lambda\mathbf{I}_{n})^{-1}\mathbf{Y}=\mathcal{K}\left(\mathbf{x},\mathbf{X}\right)\boldsymbol{\alpha},$ (12)
@@ Appendix C Proof of Theorems
$\boldsymbol{\alpha}^{\left(n\right)}=\left(\mathcal{K}\left(\mathbf{X}_{e}^{\left(1:n\right)},\mathbf{X}_{e}^{\left(1:n\right)}\right)+\lambda\mathbf{I}\right)^{-1}\mathbf{Y}^{\left(1:n\right)}.$ (21)
@@
The kernel matrix at $n$-th step for $\boldsymbol{\alpha}^{\left(n\right)}$ can be partitioned as:
@@
where the matrix $\mathbf{Y}^{\left(n\right)}$ consists of one-hot labels that are disjoint with those in previous $n-1$ domains. In this way, the parameter $\boldsymbol{\alpha}^{\left(n\right)}$ is solved by updating $\mathbf{K}^{\left(n\right)}$ and $\mathbf{C}^{\left(n\right)}$, resulting in an identical solution to the one of joint learning in Eqn. 21. Thus, the Theorem 2 is proved.